simplify the rational expression x^2+6x-16/x^2+18x+80
a. x-2/x+10
b. x-10/x+2
c. x+2/x-10
d. x+10/x-2

Answer by malakumar_kos@yahoo.com(315)


simplify the rational expression x^2+6x-16/x^2+18x+80
a. x-2/x+10
b. x-10/x+2
c. x+2/x-10
d. x+10/x-2

Factorising : x^2+6x-16 = x^2+8x-2x-16 = x(x+8)-2(x+8) = (x-2)(x+8)

x^2+18x+80 = x^2+10x+8x+80 = x(x+10)+8(x+10)
= (x+10).(x+8)
x^2+6x-16/x^2+18x+80 = (x-2).(x+8)/(x+10).(x+8)
= (x-2)/(x+10)